# Rigidity of homogeneous Lamé systems

### v1

@@ 1. Uniqueness of Lamé triplets
Theorem 1.1 is a corollary of Theorem 1.2 stated below because the homogeneous model satisfies the geometric assumptions. We next define more carefully the concepts appearing in these theorems.
Let $\Omega\subset\mathbb{R}^{3}$ be a smooth, bounded, and strictly convex domain with smooth boundary $\partial\Omega$. We refer to the Lamé parameters $(\lambda,\mu)$ and the density $\rho$ together as a Lamé triplet. We assume the functions $\lambda,\mu,\rho\colon\overline{\Omega}\to\mathbb{R}$ to be smooth and satisfy positivity conditions we will lay out in a moment.
A Lamé triplet $(\lambda,\mu,\rho)$ gives rise to an isotropic stiffness tensor field
$c_{ijk\ell}=\lambda\delta_{ij}\delta_{k\ell}+\mu\left(\delta_{ik}\delta_{j\ell}+\delta_{i\ell}\delta_{jk}\right)$ (1.2)
that is fully symmetric and depends only on the two Lamé parameters $(\lambda,\mu)$ instead of 21 independent parameters of an anisotropic stiffness tensor field, which is required to only satisfy the minimal symmetry conditions:
@@ A Lamé triplet $(\lambda,\mu,\rho)$ gives rise to an isotropic stiffness tensor field
The corresponding elastic wave operator now reads
@@
For the elastic wave operator to be hyperbolic, we require that the density $\rho$ is positive and that the spatial operator in equation (1.4) is elliptic. This, as it turns out [^1] This ellipticity amounts to requiring that the $n$-dimensional stiffness tensor $c$ is positive definite as a quadratic form on the space of symmetric real $n\times n$ matrices. That is, $c(A,A)\coloneqq A_{ij}c_{ijk\ell}A_{k\ell}>0$ for all symmetric non-zero $A\in\mathbb{R}^{n\times n}$. First, take a matrix for which $A_{12}=A_{21}=1$ and all other entries vanish. The requirement in this case is equivalent to $\mu>0$. Let us then consider a general $A$. Using the Cauchy–Schwarz inequality on the vector of eigenvalues of $A$ shows that $\operatorname{tr}(A)^{2}\leq n\operatorname{tr}(A^{2})$. Using the Hilbert–Schmidt norm, we thus get $c(A,A)=\lambda\operatorname{tr}(A)^{2}+2\mu\operatorname{tr}(A^{2})\geq(n\lambda+2\mu)\left\lvert A\right\rvert^{2}$, with equality when $A$ is a multiple of the identity. This gives the other condition $n\lambda+2\mu>0$. In the trivial case $n=1$, the condition $\mu>0$ becomes unnecessary, but the condition $\lambda+2\mu>0$ remains correct. , is equivalent to
@@
We would like to mention that the existence of a strictly convex function does not imply simplicity, as the former condition allows the existence of conjugate points while the latter forbids it. It is a major open problem whether simple manifolds admit a strictly convex function in dimension $n\geq 3$. An affirmative answer to this question would lead to major breakthroughs in geometric inverse problems. We address this further in the literature review after we have stated our main result.
@@
    the first Lamé triplet $(\lambda_{1},\mu_{1},\rho_{1})$ satisfies (1.5) and the resulting $p$- and $s$-geometries $g_{1,p/s}$ of $\Omega$ are simple and $g_{1,s}$ admits a smooth strictly convex function in $\Omega$;
@@
    the second Lamé triplet $(\lambda_{2},\mu_{2},\rho_{2})$ satisfies (1.5);
@@
    $T>\operatorname{diam}_{g_{1,s}}(\Omega)$.
@@
Remark 1.1 (Novelty of this paper).
In comparison to earlier works, in Theorem 1.2 we set very strong conditions for the first Lamé triplet, while we require nothing but the necessary positivity (1.5) from the second triplet. Since the first pair of metrics $(g_{1,p},g_{1,s})$ is both simple and admits a smooth strictly convex function, the proof of our main theorem combines results and methods presented in [11, 35, 36, 47] by building on both of these geometric conditions. It is not known to us that only one of these assumptions would be sufficient by itself. In particular, we are not aware of a counterpart of Lemma 2.4 for the convex foliation condition, where one assumes that only one of the metrics satisfies this condition. For instance, in [40, Theorem 1.3.] it is assumed that both metrics admit a strictly convex function. Neither are we cognizant of the result, such as Lemma 2.5 for simple $s$-wave (or $p$-wave) speeds. To emphasize, in the simple analog of Lemma 2.5, presented [37], the equation $\rho_{1}=\rho_{2}$ is obtained outside the locus $\lambda_{i}=2\mu_{i}$ under the assumption that the $p$-wave sectional curvature can be at most slightly positive.
Remark 1.2.
The conclusion of Theorem 1.2 remains true if, in condition (b), instead of assuming that $g_{1,s}$ admits a strictly convex function, we assume that $g_{2,s}$ has this property. This holds since we prove that $c_{1,s}=c_{2,s}$ before using the foliation condition.
@@
Since the days of [19], the study of the boundary rigidity problem has seen many breakthroughs. It was conjectured in [29] that all simple Riemannian manifolds are boundary rigid. In two dimensions, this was verified in [34]. In higher dimensions, the boundary rigidity problem is still open. However, the boundary distance rigidity holds in a conformal class of simple metrics [11], while a local version of this result was presented in [40]. Metrics near Euclidean are boundary rigid [10], a generic simple metric is boundary rigid [39]. In their groundbreaking paper [42], Stefanov–Uhlmann–Vasy prove that any three or higher-dimensional Riemannian manifold that admits a strictly convex function is boundary rigid. It is not known if all simple $n$-manifolds for $n\geq 3$ satisfy the strictly convex foliation condition. This is a major open problem in geometric inverse problems since establishing this would, in combination with [42], settle Michelle’s conjecture [29]. We would also like to mention the recent survey [43], and the excellent textbook [31] on this topic.
Very recently, Oksanen–Rakesh–Salo proved in [30] the analogue rigidity result for the Lorentzian Calderón problem. In this work, the authors showed that if a globally hyperbolic Lorentzian metric in $\mathbb{R}^{n+1}$ agrees with the Minkowski metric outside a compact set and has the same DN-map as the Minkowski metric then this metric agrees with the Minkowski metric modulo some diffeomorphism of $\mathbb{R}^{n+1}$ that agrees with the identity outside a larger compact set. Again, in this paper, no additional assumptions are imposed for the second metric. We would like to point out that in the elliptic case, the Calderón problem — which asks whether the elliptic DN-map determines a Riemannian manifold — such a rigidity result is not known. Actually, even the local version, where the second metric is assumed to be close to the Euclidean metric, remains unsolved.
@@
In comparison to our main result, the papers [41, 47] show that the Lamé triplets agree whenever their DN-maps are the same and the convex foliation condition holds for both pairs of metrics $(g_{p,i},g_{s,i})$ and $i\in\{1,2\}$.
@@ 2. Proofs
Therefore $S=\hat{S}$. The second fundamental form is $(V,W)\mapsto g(V,SW)$, so this is also invariant under the kind of conformal change we have. ∎
@@
$d_{i,p/s}(z,w)=\min\{\tau_{\hat{z}z}^{i,p/s}-\delta_{\hat{z}z}^{p/s};e_{\hat{z}z}^{i,p/s}=w,\hat{z}\text{ is as above}\}.$ (2.10)
@@
A compact Riemannian manifold $(M,g)$ with strictly convex boundary is simple if and only if the boundary distance function $d_{g}|_{\partial M\times\partial M}$ is smooth away from the diagonal [32, Proposition 3.8.14]. We are working with conformally Euclidean metrics that enjoy the rigidity of the following lemma, so we do not need this characterization of simplicity in our proof.

### v2

@@ 1. Uniqueness of Lamé triplets
Theorem 1.1 is a corollary of Theorem 1.2 stated below because the homogeneous model satisfies the geometric assumptions. We next define more carefully the concepts appearing in these theorems. Let $\Omega\subset\mathbb{R}^{3}$ be a smooth, bounded, and strictly convex domain with smooth boundary $\partial\Omega$. We refer to the Lamé parameters $(\lambda,\mu)$ and the density $\rho$ together as a Lamé triplet. We assume the functions $\lambda,\mu,\rho\colon\overline{\Omega}\to\mathbb{R}$ to be smooth and satisfy positivity conditions we will lay out in a moment.
A Lamé triplet $(\lambda,\mu,\rho)$ gives rise to an isotropic stiffness tensor field defined pointwise by
$c_{ijk\ell}=\lambda\delta_{ij}\delta_{k\ell}+\mu\left(\delta_{ik}\delta_{j\ell}+\delta_{i\ell}\delta_{jk}\right).$ (1.2)
Stiffness tensors of this special type depend only on the two Lamé parameters $(\lambda,\mu)$ instead of 21 independent parameters of an anisotropic stiffness tensor field, which is required to only satisfy the minimal symmetry conditions
@@
The elastic wave operator $P$ is defined by
@@
For the elastic wave operator to be hyperbolic, we require that the density $\rho$ is positive and that the spatial operator in equation (1.4) is elliptic. This, as it turns out [^1] This ellipticity amounts to requiring that the $n$-dimensional stiffness tensor $c$ is positive definite as a quadratic form on the space of symmetric real $n\times n$ matrices. That is, $c(A,A)\coloneqq A_{ij}c_{ijk\ell}A_{k\ell}>0$ for all symmetric non-zero $A\in\mathbb{R}^{n\times n}$. First, take a matrix for which $A_{12}=A_{21}=1$ and all other entries vanish. The requirement in this case is equivalent to $\mu>0$. Let us then consider a general $A$. Using the Cauchy–Schwarz inequality on the vector of eigenvalues of $A$ shows that $\operatorname{tr}(A)^{2}\leq n\operatorname{tr}(A^{2})$. Using the Hilbert–Schmidt norm, we thus get $c(A,A)=\lambda\operatorname{tr}(A)^{2}+2\mu\operatorname{tr}(A^{2})\geq(n\lambda+2\mu)\left\lvert A\right\rvert^{2}$, with equality when $A$ is a multiple of the identity. This gives the other condition $n\lambda+2\mu>0$. In the case $n=1$, the condition $\mu>0$ becomes unnecessary, but the condition $\lambda+2\mu>0$ remains correct. , is equivalent to
@@
The existence of a strictly convex function does not imply simplicity, as the former condition allows the existence of conjugate points while the latter forbids it. It is an open problem whether simple manifolds admit a strictly convex function in dimension $n\geq 3$.
@@
    $T>\operatorname{diam}_{g_{1,s}}(\Omega)$ in the sense of the distance induced by $g_{1,s}$ on $\overline{\Omega}$.
@@
    Both Lamé triplets $(\lambda_{i},\mu_{i},\rho_{i})$ satisfy (1.5).
@@
    $(\overline{\Omega},g_{1,p})$ is a simple Riemannian manifold.
  • (e)
    $(\overline{\Omega},g_{1,s})$ is simple and admits a strictly convex foliation.
@@
A few points about the nature of our assumptions is in order:
  • Asymmetric assumptions between models:
    • We make stringent geometric assumptions on model 1 (i.e. the Lamé triplet $(\lambda_{1},\mu_{1},\rho_{1})$), while no geometric assumptions whatsoever are made on model 2.
    • Similar asymmetric rigidity results exist in the literature. In [40, Theorem 1.4.] only one of the two metrics compared is assumed to admit a strictly convex foliation. In [11] (which we use as lemma 2.5) and [47] only one of the two compared Riemannian metrics is assumed to be simple.
  • Asymmetric assumptions between polarizations: Our geometric assumptions are different for $p$ and $s$ wave metrics in model 1.
  • Tradeoff between measurement time and geometric assumptions:
    • Our measurement time $T$ is only assumed to exceed the $g_{1,s}$-diameter of the domain in lemma 2.2, not the lengths of all geodesics. Once we assume simplicity in theorem 1.2, these two quantities agree.
    • Our reduction from PDE data to geometric data produces boundary distance data, not lens data (i.e. scattering relation and lengths of geodesics; see e.g. [40]). Lens data is compatible with the foliation condition while boundary distance data is not, which is why we make the additional assumption of simplicity of $g_{1,p/s}$.
    • It is possible to formulate a similar result to theorem 1.2 with foliation conditions only on the metrics $g_{1,p/s}$. The assumptions differ from ours in a few ways: The measurement time is longer than all $g_{1,s}$-geodesics, not just greater than the $g_{1,s}$-diameter. Both metrics $g_{1,p/s}$ are assumed to have a strictly convex foliation, $\partial\Omega$ is strictly convex with respect to $g_{1,p/s}$, and neither metric is assumed simple. Here is a sketchy argument to support this: The overall idea is similar to the one we give for our main results in this paper, making use of [35], [40, Theorem 1.4.], [41, Theorems 2], and [48, Theorem 1]. However, in order to recover the lens data instead of the boundary distance function, one needs to replace our lemma 2.2 with [41, Theorems 2]. To our best knowledge, this result has not appeared in the literature yet.
@@
Since the days of [19], the study of the boundary rigidity problem has seen many breakthroughs. It was conjectured in [29] that all simple Riemannian manifolds are boundary rigid. In two dimensions, this was verified in [34]. In higher dimensions, the boundary rigidity problem is still open. However, the boundary distance rigidity holds in a conformal class of simple metrics [11], while a local version of this result was presented in [40]. Metrics near Euclidean are boundary rigid [10], a generic simple metric is boundary rigid [39]. In their groundbreaking paper [42], Stefanov–Uhlmann–Vasy prove that any three or higher-dimensional Riemannian manifold that admits a strictly convex function is boundary rigid. It is not known if all simple $n$-manifolds for $n\geq 3$ satisfy the strictly convex foliation condition. This is a major open problem in geometric inverse problems since establishing this would, in combination with [42], settle Michel’s conjecture [29]. We would also like to mention the recent survey [43], and the excellent textbook [31] on this topic.
Very recently, Oksanen–Rakesh–Salo proved in [30] the analogues rigidity result for the Lorentzian Calderón problem. In this work, the authors showed that if a globally hyperbolic Lorentzian metric in $\mathbb{R}^{n+1}$ agrees with the Minkowski metric outside a compact set and has the same DN-map as the Minkowski metric then this metric agrees with the Minkowski metric modulo some diffeomorphism of $\mathbb{R}^{n+1}$ that agrees with the identity outside a larger compact set. Again, in this paper, no additional assumptions are imposed for the second metric. We would like to point out that in the elliptic case, the Calderón problem — which asks whether the elliptic DN-map determines a Riemannian manifold — such a rigidity result is not known. Actually, even the local version, where the second metric is assumed to be close to the Euclidean metric, remains unsolved.
@@ 2. Proofs
Therefore the shape operators of $g$ and $\hat{g}$ satisfy $S=\hat{S}$. The second fundamental form is $(V,W)\mapsto g(V,SW)$, so this is also invariant under the kind of conformal change we have. ∎
@@
$d_{i,p/s}(z,w)=\min\{\tau_{\hat{z}z}^{i,p/s}-\delta_{\hat{z}z}^{p/s};e_{\hat{z}z}^{i,p/s}=w\text{ and }\hat{z}\text{ is as above}\}.$ (2.10)
@@
A compact Riemannian manifold $(M,g)$ with strictly convex boundary is simple if and only if the boundary distance function $d_{g}|_{\partial M\times\partial M}$ is smooth away from the diagonal [32, Proposition 3.8.14]. We are working with conformally Euclidean metrics that enjoy the rigidity of the following lemma, so we do not directly need this characterization of simplicity in our proof.